# Unified Manipulability and Compliance Analysis of Modular Soft-Rigid Hybrid Fingers

### v1

@@ 1 Introduction
Achieving dexterous and robust manipulation in unstructured environments remains a core challenge in robotics Billard and Kragic (2019). As robotic hands advance beyond simple grasping toward versatile in-hand manipulation, there is growing demand for systems that combine precise control, adaptability, compliance, and safe physical interaction. Soft robotic systems excel in adaptive grasping through material compliance and morphology-driven control but face limitations in precision tasks. Soft-rigid hybrid systems, which integrate compliant actuators with rigid linkages, offer a balanced approach—enabling both fine control and structural adaptability.
@@
A variety of soft and soft-rigid robotic hands has emerged over the past decade. Pneumatic systems like the RBO Hand Deimel and Brock (2016) and the Pisa/IIT SoftHand Catalano et al. (2014) enable compliant grasping via soft inflation or tendon-driven underactuation. Hybrid hands enhance control and robustness by combining soft actuators with rigid structures, including origami mechanisms Li et al. (2019), layer jamming Manti et al. (2016), bellow-type encapsulation Lee et al. (2024), and geometric confinement Zhang et al. (2020). Variable stiffness Wolf et al. (2015) and programmable actuation Polygerinos et al. (2017) further boost adaptability.
@@
Our prior work explores both soft and soft-rigid hands. The gripper in Zhou et al. (2017) uses textured dual-chamber fingers for low-pressure adaptation. The BCL-26 hand Zhou et al. (2019) achieves human-like motion with 26 fiber-reinforced soft DOFs. We incorporated stiffness tuning via particle jamming in Zhou et al. (2020). The DexCo Hand Zhou et al. (2024) uses hydraulic actuation for forceful, controllable motion, while the Edgy-2 Hand Zhou et al. (2018) emphasizes modular pneumatic flexibility.
@@
Despite recent advances, many studies remain focused on empirical evaluation and lack generalizable models for analyzing or comparing design variants. While task-specific tuning may suffice for grasping, manipulation tasks demand predictive insights into how structure and actuation affect performance. Two key metrics—manipulability, describing motion capability Doty et al. (1995), and compliance, capturing passive deformation under force Aukes et al. (2014)—are essential for assessing dexterity and adaptability. Though common in rigid and underactuated systems, these metrics are rarely jointly analyzed in soft-rigid hybrid designs, especially across different actuation modes. This gap hinders structure-aware design and cross-system performance generalization.
@@
To address this, we propose a unified analytical framework for evaluating both manipulability and compliance in modular soft-rigid hybrid fingers actuated by hydraulic or pneumatic systems. We formulate a Jacobian-based model that maps actuator inputs to joint and task-space behaviors. Hydraulic systems are modeled using a linear approximation under incompressible flow, while pneumatic actuators are captured using nonlinear pressure–volume relationships derived from ideal gas behavior Polygerinos et al. (2017). This formulation enables consistent computation of manipulability ellipsoids and compliance matrices across actuation types. We validate the proposed framework on two representative systems: the hydraulically actuated DexCo Hand and the pneumatically actuated Edgy-2 Hand. Simulation and experimental results reveal actuation-dependent trade-offs in directional dexterity and passive stiffness, offering insights for structure-aware analysis, actuator selection, and modular finger design in soft-rigid robotic fingers.
@@
Figure 1: (a) Visualization of the model that approximates linear soft actuators. Dashed lines labeled with J1, J2 and J3 marks the positions of the passive revolute joints. Solid lines labeled with L1, L2 and L3 are the linear soft actuator. (b)The top down view of one of the fingers of the Dexco hand when $\theta_{1}=0$. The parameters used in (1) and (2) are labeled.
@@ 2.1 Dual Soft Actuation Case
Here, $A$, $H$, and $V$ represent the axial, horizontal, and vertical offsets, respectively, from joint $J_{1}$ to joint $J_{2}$, as functions of $\theta_{2}$ when $\theta_{1}=0$. These offsets can be computed as follows. The parameter $R$ denotes the radial distance from the attachment points of actuators $L_{1}$ and $L_{2}$ to the axis of joint $J_{2}$. The quantities $D_{1}$ and $D_{2}$ represent the offsets between the actuator attachment surface and the rotation axes. Fig. 1b marks a visualization of these parameters. Lastly, $L$ is the distance between the actuator endpoints on the attachment surface.
@@
To further analyze the relationship between actuator inputs and joint configurations, we plot the projected curves along the contour directions, as shown in Fig. 3a. The two contour paths are defined from point A to B and from point C to D, as indicated in Fig. 2. The x-axis values in these plots correspond to a normalized actuator motion (denoted as $NM$). Taking the contour from A to B as an example, for a given point $(l_{1},l_{2},\theta_{1})$ on the contour, and letting $(l_{1,A},l_{2,A},\theta_{1})$ denote the coordinates at point A, the normalized actuator motion is computed as follows:
$NM=\sqrt{(l_{1}-l_{1,A})^{2}+(l_{2}-l_{2,A})^{2}}$ (3)
@@
Fig. 4 presents the Jacobian analysis over the entire workspace. The Jacobian is computed using the following equation, which characterizes the actuator effort required to produce joint motions in arbitrary directions.
$J=\left|\det{\begin{bmatrix}\frac{\partial\theta_{1}}{\partial l_{1}}&\frac{\partial\theta_{1}}{\partial l_{2}}\\ \frac{\partial\theta_{2}}{\partial l_{1}}&\frac{\partial\theta_{2}}{\partial l_{2}}\end{bmatrix}}\right|=\left(\left|\det{\begin{bmatrix}\frac{\partial l_{1}}{\partial\theta_{1}}&\frac{\partial l_{1}}{\partial\theta_{2}}\\ \frac{\partial l_{2}}{\partial\theta_{1}}&\frac{\partial l_{2}}{\partial\theta_{2}}\end{bmatrix}}\right|\right)^{-1}$ (4)
The partial derivatives are directly obtained from (1). To eliminate the influence of unit selection, the Jacobian is normalized by dividing it by its value at the reference configuration where $\theta_{1}=\theta_{2}=0$. As shown in Fig. 4, the normalized Jacobian remains well-conditioned across the workspace, without exhibiting any extreme or anomalous values.
Figure 5: Experimental setup. The joint angles are captured from ArUco code (Romero-Ramirez et al. (2018)).
@@ 2.3 Experimental Validation with Hydraulic Actuators
Due to the near-incompressibility of the liquid under room temperature and moderate pressure, the actuator lengths $l$ are expected to maintain an affine relationship with the piston positions, i.e., $l=kQ+b$, where $k$ and $b$ are scalar coefficients determined experimentally. Based on this assumption, we collected 200 data samples of the form $(Q_{1}^{e},Q_{2}^{e},Q_{3}^{e},\theta_{1}^{e},\theta_{2}^{e},\theta_{3}^{e})$, and computed the corresponding expected actuator lengths $l_{1}^{e}$, $l_{2}^{e}$, and $l_{3}^{e}$. Linear regression was then performed between the pairs $(Q_{1}^{e},l_{1}^{e})$, $(Q_{2}^{e},l_{2}^{e})$, and $(Q_{3}^{e},l_{3}^{e})$, as illustrated in Fig. 6. The results demonstrate a strong linear correlation between actuator lengths and piston positions, with high coefficients of determination, validating the geometric models in (1) and (5).
@@ 2.4 Potential Extension to Pneumatic Actuators
Figure 7: Illustration of compliance modeling during the force estimation stage in actuation space.
@@ 3 Compliance Modeling
While manipulability characterizes the motion capability of the fingers, compliance captures how they respond to external forces—an equally critical factor in soft-rigid design. Due to the inherent compliance of soft materials, both hydraulic and pneumatic actuators exhibit compliant behavior during interaction. In this section, we focus on the static analysis and compliance modeling of the DexCo hand, mapping from the actuation space ($\boldsymbol{a}$) to the configuration space ($\boldsymbol{\tau}$). We develop models to characterize the compliance introduced by both hydraulic and pneumatic actuation, and validate these models through simulations involving object manipulation.
@@ 3.1 Hydraulic Compliance Modeling
In hydraulic systems, compliance arises primarily from the compressibility of the working fluid, elasticity of the fluidic chamber and tubing, and mechanical compliance of the structure. Although hydraulic actuation is often regarded as high-stiffness and low-compliance, for soft-rigid hybrid systems, even small compliance effects can significantly influence control and force interaction performance.
The hydraulic force in soft actuators arises from the elastic deformation of the material in the radial direction during compression or extension, similar to inflating a balloon (Fig. 7).
@@
$\varepsilon_{1}=\frac{\sigma_{1}}{E}=\frac{\Delta_{1}}{2\pi r}$ (8)
where $\Delta_{1}$ is the change in circumference due to pressure. Using (7) and (8), we derive the expressions for the deformed inner and outer diameters under pressure $P$:
@@ 3.3 Simulate Compliance Motion
$M(\theta)\,\ddot{\theta}+H(\dot{\theta},\theta)=\tau_{k}$ (17)
In Eq. 17, the left-hand side represents the inertial, Coriolis, and gravitational forces, which are computed by the physics simulator (gazebo). The right-hand side, $\tau_{k}$, denotes the torque induced by elastic deformation. By incorporating $\tau_{k}$ into the simulation, we are able to replicate the effects of nonlinear, variable stiffness elastic forces on the hand’s behavior, as shown in the Fig. 8B.
In the implemented simulation, we conducted a motion sequence consisting of two phases. First, a compliant caging grasp was performed to securely enclose the object. Then, by increasing the actuation force at the base joint of the left finger while altering the orientation of the fingertip, the object was deliberately pushed away. Figure 8A illustrates the commanded joint angles alongside the actual joint responses. Due to the variable compliance of the system, discrepancies arise between the desired and actual joint angles. Joints 2 and 3 correspond to the universal joint at the finger base, while Joint 4 represents the fingertip joint.
Figure 8: Simulation in gazebo of the compliant motions of the DexCo hand.
@@ 4 Conclusion
This paper proposed a unified framework for analyzing the manipulability and compliance of modular soft-rigid hybrid fingers actuated by hydraulic and pneumatic systems. A generalized Jacobian-based formulation was introduced to model the actuator-to-task-space relationship, enabling consistent evaluation of directional dexterity and passive compliance across actuation types. Simulations and experiments on the DexCo and Edgy-2 fingers validated the framework and highlighted key trade-offs between actuation modality, structural configuration, and control sensitivity.
The results provide a foundation for structure-aware and actuation-aware design of soft-rigid fingers, facilitating informed selection and optimization in modular robotic systems. Future work will extend this framework to include dynamic effects such as valve delay and flow regulation in pneumatic actuation, as well as integrate contact modeling for task-level compliance control and grasp planning.

### v2

@@ 1 Introduction
Achieving dexterous and robust manipulation in unstructured environments remains a core challenge in robotics (Billard and Kragic, 2019). As robotic hands advance beyond simple grasping toward versatile in-hand manipulation, there is growing demand for systems that combine precise control, adaptability, compliance, and safe physical interaction. Soft robotic systems excel in adaptive grasping through material compliance and morphology-driven control but face limitations in precision tasks. Soft-rigid hybrid systems, which integrate compliant actuators with rigid linkages, offer a balanced approach—enabling both fine control and structural adaptability.
@@
A variety of soft and soft-rigid robotic hands has emerged over the past decade. Pneumatic systems like the RBO Hand (Deimel and Brock, 2016) and the Pisa/IIT SoftHand (Catalano et al., 2014) enable compliant grasping via soft inflation or tendon-driven underactuation. Hybrid hands enhance control and robustness by combining soft actuators with rigid structures, including origami mechanisms (Li et al., 2019), layer jamming (Manti et al., 2016), bellow-type encapsulation (Lee et al., 2024), and geometric confinement (Zhang et al., 2020). Variable stiffness (Wolf et al., 2015) and programmable actuation (Polygerinos et al., 2017) further boost adaptability.
@@
Our prior work explores both soft and soft-rigid hands. The gripper in (Zhou et al., 2017) uses textured dual-chamber fingers for low-pressure adaptation. The BCL-26 hand (Zhou et al., 2019) achieves human-like motion with 26 fiber-reinforced soft degrees of freedom (DOFs). We incorporated stiffness tuning via particle jamming in (Zhou et al., 2020). The DexCo Hand (Zhou et al., 2024) uses hydraulic actuation for forceful, controllable motion, while the Edgy-2 Hand (Zhou et al., 2018) emphasizes modular pneumatic flexibility.
@@
Despite recent advances, many studies remain focused on empirical evaluation and lack generalizable models for analyzing or comparing design variants. While task-specific tuning may suffice for grasping, manipulation tasks demand predictive insights into how structure and actuation affect performance. Two key metrics—manipulability, describing motion capability (Doty et al., 1995), and compliance, capturing passive deformation under force (Aukes et al., 2014)—are essential for assessing dexterity and adaptability. Though common in rigid and underactuated systems, these metrics are rarely jointly analyzed in soft-rigid hybrid designs, especially across different actuation modes. This gap hinders structure-aware design and cross-system performance generalization.
@@
To address this, we propose a unified analytical framework for evaluating both manipulability and compliance in modular soft-rigid hybrid fingers actuated by hydraulic or pneumatic systems. We formulate a Jacobian-based model that maps actuator inputs to joint and task-space behaviors. Hydraulic systems are modeled using a linear approximation under incompressible flow, while pneumatic actuators are captured using nonlinear pressure–volume relationships derived from ideal gas behavior (Polygerinos et al., 2017). This formulation enables consistent computation of manipulability ellipsoids and compliance matrices across actuation types. We validate the proposed framework on two representative systems: the hydraulically actuated DexCo Hand and the pneumatically actuated Edgy-2 Hand. Simulation and experimental results reveal actuation-dependent trade-offs in directional dexterity and passive stiffness, offering insights for structure-aware analysis, actuator selection, and modular finger design in soft-rigid robotic fingers.
@@
Figure 1: (a) Kinematic model of linear soft actuators: dashed lines (J1–J3) denote passive revolute joints; solid lines (L1–L3) represent soft actuators. (b) Top view of a DexCo finger at $\theta_{1}=0$, with parameters used in (1) and (2) labeled.
@@ 2.1 Dual Soft Actuation Case
Here, $A$, $H$, and $V$ represent the axial, horizontal, and vertical offsets, respectively, from joint $J_{1}$ to joint $J_{2}$, as functions of $\theta_{2}$ when $\theta_{1}=0$. These offsets can be computed as follows. The parameter $R$ denotes the radial distance from the attachment points of actuators $L_{1}$ and $L_{2}$ to the axis of joint $J_{2}$. The quantities $D_{1}$ and $D_{2}$ represent the offsets between the actuator attachment surface and the rotation axes. The parameter $S$ denotes the horizontal distance from joint $J_{1}$ to the actuator base attachment point, as shown in Fig. 1b. Lastly, $L$ is the distance between the actuator endpoints on the attachment surface.
@@
To further analyze the relationship between actuator inputs and joint configurations, we plot the projected curves along the contour directions, as shown in Fig. 3a. The two contour paths are defined from point A to B and from point C to D, as indicated in Fig. 2. The x-axis values in these plots correspond to a normalized actuator motion (denoted as $NAM$). Taking the contour from A to B as an example, for a given point $(l_{1},l_{2},\theta_{1})$ on the contour, and letting $(l_{1,A},l_{2,A},\theta_{1})$ denote the coordinates at point A, the normalized actuator motion is computed as follows:
$NAM=\sqrt{(l_{1}-l_{1,A})^{2}+(l_{2}-l_{2,A})^{2}}$ (3)
@@
Fig. 4 presents the manipulability, $M$, analysis over the entire workspace. The manipulability is computed using the following equation, which characterizes the actuator effort required to produce joint motions in arbitrary directions.
$M=\left|\det{\begin{bmatrix}\frac{\partial\theta_{1}}{\partial l_{1}}&\frac{\partial\theta_{1}}{\partial l_{2}}\\ \frac{\partial\theta_{2}}{\partial l_{1}}&\frac{\partial\theta_{2}}{\partial l_{2}}\end{bmatrix}}\right|=\left(\left|\det{\begin{bmatrix}\frac{\partial l_{1}}{\partial\theta_{1}}&\frac{\partial l_{1}}{\partial\theta_{2}}\\ \frac{\partial l_{2}}{\partial\theta_{1}}&\frac{\partial l_{2}}{\partial\theta_{2}}\end{bmatrix}}\right|\right)^{-1}$ (4)
The partial derivatives are directly obtained from (1). To eliminate the influence of unit selection, the manipulability can be normalized by dividing it by its value at the reference configuration where $\theta_{1}=\theta_{2}=0$. As shown in Fig. 4, the normalized manipulability remains well-conditioned across the workspace, without exhibiting any extreme or anomalous values.
Figure 5: Experimental setup. The joint angles are captured from ArUco code ((Romero-Ramirez et al., 2018)).
@@ 2.3 Experimental Validation with Hydraulic Actuators
Due to the near-incompressibility of the liquid under room temperature and moderate pressure, the actuator lengths $l$ are expected to maintain an affine relationship with the piston positions, i.e., $l=kQ+b$, where $k$ and $b$ are scalar coefficients determined experimentally. Based on this assumption, we collected 200 data samples of the form $(Q_{1}^{e},Q_{2}^{e},Q_{3}^{e},\theta_{1}^{e},\theta_{2}^{e},\theta_{3}^{e})$, and computed the corresponding expected actuator lengths $l_{1}^{e}$, $l_{2}^{e}$, and $l_{3}^{e}$. Linear regression was then performed between the pairs $(Q_{1}^{e},l_{1}^{e})$, $(Q_{2}^{e},l_{2}^{e})$, and $(Q_{3}^{e},l_{3}^{e})$, as illustrated in Fig. 6. The results demonstrate a strong linear correlation between actuator lengths and piston positions, with high coefficients of determination, validating the geometric models in (1) and (5). The slightly higher variance observed in the L1 data may be attributed to minor misalignment or increased friction in its tubing connection, which affects the repeatability of piston-to-length translation in the physical setup.
@@ 2.4 Potential Extension to Pneumatic Actuators
Figure 7: (A) Compliance modeling for force estimation in actuation space. (B) Initial state of the origami bellow. (C) Extended and contracted states.
@@ 3 Compliance Modeling
While manipulability characterizes motion capability, compliance captures the response to external forces—an equally critical aspect in soft-rigid design. Owing to material softness, both hydraulic and pneumatic actuators exhibit inherent compliance during interaction. This section presents a static analysis and compliance modeling of the soft-rigid hybrid hand, mapping actuation inputs ($\boldsymbol{a}$) to joint configurations ($\boldsymbol{\tau}$). We develop models for the compliance introduced by hydraulic and pneumatic actuation and validate them through object manipulation simulations.
@@ 3.1 Hydraulic Compliance Modeling
In hydraulic systems, compliance arises primarily from the compressibility of the working fluid, the elasticity of the fluidic chamber and tubing, and the mechanical compliance of the structure.
Compared to traditional hydraulic systems that exhibit high stiffness and low compliance, the dominant source of compliance in soft hydraulic actuators is the elastic deformation of the chamber material, characterized by its Young’s modulus $E$. As the material deforms under pressure, the internal diameter changes from its undeformed size $D_{0}$ and $d_{0}$ to the deformed value $D$ and $d$ thereby altering the internal geometry and enabling displacement. The hydraulic force in soft actuators arises from the elastic deformation of the material in the axial direction during compression or extension, similar to inflating a balloon (Fig. 7).
@@
$\varepsilon_{1}=\frac{\sigma_{1}}{E}=\frac{\Delta_{\ell}}{2\pi r}$ (8)
where $\Delta_{\ell}$ denotes the change in circumference due to radial expansion, i.e., $\Delta_{\ell}=2\pi(r-r_{0})$, with $r_{0}$ representing the undeformed radius and $r$ the deformed radius. Using (7) and (8), we derive the expressions for the deformed inner and outer diameters under pressure $P$ via replacing $r$ with $\frac{d_{0}}{2}$ and $\frac{D_{0}}{2}$:
@@ 3.2 Pneumatic Actuation Model
Experimental validation of pressure–deformation behavior in our previous work  (Zhou et al., 2018) supports the pneumatic modeling in this work.
@@ 3.3 Simulate Compliance Motion
$M(\theta)\,\ddot{\theta}+H(\dot{\theta},\theta)+G=\tau_{k}$ (17)
In Eq. 17, the left-hand side captures inertial, Coriolis, and gravitational effects computed by the Gazebo simulator. The right-hand term, $\tau_{k}$, represents torques from elastic deformation. Incorporating $\tau_{k}$ enables simulation of nonlinear variable stiffness behaviors, as shown in Fig. 8B.
The simulation involves two phases: a compliant caging grasp to enclose the object, followed by forceful release via increased base joint actuation and fingertip reorientation. Fig. 8A shows both commanded and actual joint angles. Compliance-induced deviations are observed between target and actual responses, where Joints 2 and 3 form the base, and Joint 4 corresponds to the fingertip.
Figure 8: Compliant motion simulation of the Soft-Rigid hybrid hand in Gazebo: (A) joint trajectories; (B) final interaction state.
@@ 4 Conclusion
This paper presents a unified framework for analyzing the manipulability and compliance of modular soft-rigid hybrid fingers actuated by hydraulic and pneumatic systems. Using a generalized Jacobian-based formulation, we enable consistent structural evaluation of directional dexterity and passive compliance across actuation modes. Experimental validation on the DexCo (hydraulic) and Edgy-2 (pneumatic) hands highlights key trade-offs between actuation type, structure, and control sensitivity.
Our results highlight that control response and actuation speed are distinct in hybrid systems. As shown in Table 1, hydraulic actuators enable precise control but slower motion, while pneumatic actuators offer faster deformation with less stability. These mismatches suggest the need for functional separation or hierarchical integration in control design.
Table 1: Summary of actuation trade-offs.
Metric Hydraulic Pneumatic
Control response speed Fast Slow
Actuation speed Slower Faster
Force output precision High Low–moderate
Passive compliance Low High
We recommend structure-aware and actuation-aware design principles, including: (1) avoiding direct coupling of mismatched subsystems; (2) assigning distinct roles to each actuator type (e.g., pneumatic for grasping, hydraulic for in-hand manipulation); and (3) leveraging our framework to assess design trade-offs early in development.
While the current formulation focuses on quasi-static behavior, future work will incorporate dynamic effects such as inertia and damping, particularly for soft-rigid hybrid fingers. We also aim to extend the framework to include contact interactions and grasp stability for task-level analysis and robust control.